OpenStudy (albert0898):

If f(x) = 2 for all real numbers x, then f(x+2) = (A) 0 (B) 2 (C) 4 (D) x (E) The value cannot be determined.

OpenStudy (dannyo19):

Say we have an equation f(x)=x^2 f(2) will equal 4 while f(x+2)=4^2 or 16 Now say f(x)=x+800, f(2)=802 f(x+2)=804 problem is f(x)=2 doesn't tell us a function involving x, so how can we possibly determine f(x+2)?

OpenStudy (dannyo19):

f(x) as a function could be anything as long as f(x)=2

OpenStudy (albert0898):

The answer is B, but I perplexed as to how to get the answer. I understand that when f(x) is plugged it, 2 is yielded. Does that mean that the answer remains the same for any function?

OpenStudy (dannyo19):

But it says f(x)=2 for all real numbers x

OpenStudy (albert0898):

This book has very poor explanations, idk :O

OpenStudy (dannyo19):

So f(4)=2, f(anything)=2 agreed?

OpenStudy (dannyo19):

any real number so f(x+1000) must equal?

OpenStudy (albert0898):

2?

OpenStudy (dannyo19):

Yes

OpenStudy (albert0898):

Is there a specific rule or no?

OpenStudy (dannyo19):

So if for all real numbers f(x)=2 f(x+2) where x is a real number will be no different

OpenStudy (albert0898):

@Zarkon do you know how to solve this? I'm still very confused

OpenStudy (anonymous):

plug in 2 into x

OpenStudy (anonymous):

@Albert0898 are you with me

OpenStudy (zarkon):

"If f(x) = 2 for all real numbers x" the f(5)=2 f(y)=2 f(456436346346x+3453535)=2 whatever you plug in you get 2

OpenStudy (zarkon):

f(x)=2 has a graph that is a horizontal line with height 2

OpenStudy (zarkon):

if x is a real number then x+2 is a real number therefore f(x+2)=...

OpenStudy (anonymous):

then the answer is 4

OpenStudy (dannyo19):

f(x+2) must have an x value among all real numbers f(x) must have an x value among all real numbers (-inf,inf) So f(x)=2 Now say x=2 what will f(2) be equal to?

OpenStudy (dannyo19):

That won't change if x=4 right?

OpenStudy (dannyo19):

That won't change if f(x)=f(x+500) no matter the input x the output is 2

OpenStudy (albert0898):

I think I get it now! f(x) = y = 2 f(x+2) is equal to y which is equal to 2? @DannyO19 @Zarkon @Daniel56k

OpenStudy (danjs):

yeah the first sentence in the prob says it is same y for all x, horizontal line Y=2 f(x - a) is a way to take the whole graph of the function, and shift it horizontal 'a' units.

OpenStudy (danjs):

so after applying f(x+2) the thing shifts left 2 units this time, (a=-2) this prob is just y = constant value so really it doesnt matter, but that is that

OpenStudy (albert0898):

Oh I get it now! Thanks!

OpenStudy (danjs):

like say you choose x=2, the function will give you f(4), that is 2 units pelletf left...